OpenStudy (anonymous):

Describe how to find the percent of change in the price of a skateboard from last year to this year.

OpenStudy (dumbcow):

(newPrice - oldPrice) / oldPrice X 100%

OpenStudy (dumbcow):

Example: last year, price = $12 this year, price = $15 percent change = (15-12)/12 *100% = (3/12) *100% =(1/4)*100% = 0.25 *100% = 25%